A box contains 40 Balls, 25% of them are white whereas the rest of them are black. A person randomly
picked 4 balls from the box without replacement, find the probability of getting 3 black balls and 1
white ball.
1
Expert's answer
2020-12-06T18:50:26-0500

Given:


Total= 40

Number of white balls: 25% *40= 10

Number of black= 40-10=30


Without replacement:

P(3 black balls and 1 white ball)= "\frac{30}{40}* \frac{29}{39}*\frac{28}{38}*\frac{10}{37}"


="\frac{243600}{21,93,360}"


=0.111